Based on VSEPR theory, the number of the 90 deegre F-Br-F angles in BrF5 is

Grade:12

1 Answers

Arun
6 years ago
According to VSEPR theory, number of electron pairs around central atom (Br) are 6..N/2=(7+5)/2=6. (Five are bond pairs and one is lone pair)Its geometry is octahedral but due to lone pair- bond pair repulsion , the four fluorine atoms at corner are forced towards the upper fluorine atom thus reducing F - Br -F angle from 90 to 84.8∴∴ The answer is 0.